OpenStudy (condor34):

Simplify the expression sqrt -9/ (3-2i)+(1+5i)

Nnesha (nnesha):

first combine like terms

Nnesha (nnesha):

\[\huge\rm \frac{ \sqrt{-9} }{ 3-2i+1+5i }\]

OpenStudy (condor34):

12-9i/25 12-9i/7 9+12i/25 9+12i/7

OpenStudy (condor34):

and yes

OpenStudy (condor34):

square root of 9 over 4 plus 3i?

OpenStudy (condor34):

-9*

Nnesha (nnesha):

alright and you separate -9 \[\sqrt{-1} \times \sqrt{9}\] and i =what ?

Nnesha (nnesha):

yep right

OpenStudy (condor34):

i = imaginary number? or -1

Nnesha (nnesha):

i imaginary

Nnesha (nnesha):

general question

OpenStudy (condor34):

ok

Nnesha (nnesha):

\[\bf i = \sqrt{-1}\] \[\bf i^2 = -1\] and that's why we have to seprate of negative one and positive 9

OpenStudy (condor34):

ok

Nnesha (nnesha):

\[\huge\rm \frac{ \sqrt{-1} \times \sqrt{9} }{ 4-3i }\] so replace sqrt -1 by 1 and multiply both the denominator and the numerator by the conjugate of 4-3i

OpenStudy (condor34):

-9/25+12i/25?

Nnesha (nnesha):

nope how did you get 25

OpenStudy (condor34):

just guessed

Nnesha (nnesha):

why did you guess ?

Nnesha (nnesha):

you are not 'n the exam haall this is the time to learn not to guess :=)

OpenStudy (condor34):

Honestly I have no idea how to do this and not trying to be rude but i need the answer because i am taking a timed exam.......

Nnesha (nnesha):

really so it's a test question ..sorryy..

OpenStudy (condor34):

yes

Nnesha (nnesha):

do you know we are not allowed to help u on your test or quiz? it's against the coc

OpenStudy (condor34):

I know but i really dont understand this and i have also seen alot of people do the same thing as me

Nnesha (nnesha):

well..if someone don't aa bad thing then.. you don't have to follow them you will do better without cheating..

OpenStudy (condor34):

ok, thanks for the help anway

Nnesha (nnesha):

what is square root of 9? replace square root of -1 by i multiply by the conjugate 4-3i if you don't know what conjugate then search it